OpenStudy (anonymous):

Which isotopes have the same number of neutrons? superscript 12 over subscript 6 C superscript 19 over subscript 9 F superscript 20 over subscript 10 Ne superscript 23 over subscript 11 Na

OpenStudy (anonymous):

@NoDino

OpenStudy (anonymous):

@agreene

OpenStudy (agreene):

So, to find the number of neutrons you will want to subtract the mass number from the number of protons (protons = electrons = bottom number) 12-6 = 6 neutrons 19-9 = 10 neutrons 20 - 10 = 10 neutrons 23 - 11 = 12 neutrons so F-19 and Ne-20 have the same number of neutrons

OpenStudy (agreene):

I didn't say that very clear: Neutrons = Mass number - Protons Protons = Electrons = Atomic Number

OpenStudy (anonymous):

okay thanks for the help